OpenStudy (anonymous):

Just something to work on, Prove \( \large \sum \limits_{k=1}^{n} k\times k! = (n+1)! -1 \) without using induction.

OpenStudy (zarkon):

easy :)...use the same 'trick' I did for the last problem

OpenStudy (anonymous):

Only few things could be somewhat hard for you Zarkon :D but still not all of us know how to read your mind :)

OpenStudy (zarkon):

k=(k+1)-1

OpenStudy (anonymous):

Yes, telescopic summation :)

OpenStudy (mr.math):

I've been trying a different approach, but it doesn't seem to be working :-(

myininaya (myininaya):

ffm, zarkon has an alien brain

OpenStudy (zarkon):

that is ture

OpenStudy (zarkon):

I feel sorry for all you HUmons

OpenStudy (anonymous):

Yes, all these professors seems to have that :P

OpenStudy (asnaseer):

\[\begin{align} \sum \limits_{k=1}^{n} k\times k!&=\sum \limits_{k=1}^{n} ((k+1)-1)\times k!\\ &=\sum \limits_{k=1}^{n} (k+1)! - k!\\ &=2!-1!+3!-2!+4!-3!+...+(n+1)!-n!\\ &=(n+1)!-1 \end{align}\]using hint given by Zarkon and telescopic sum as ffm indicated

myininaya (myininaya):

I like that :)

OpenStudy (asnaseer):

all credit and trademarks belong to Zarkon and ffm :D

myininaya (myininaya):

how much do they get paid, ffm?

OpenStudy (asnaseer):

Zarkon especially - he seems to have a brain like a laser! :D

OpenStudy (zarkon):

Its Planet doom university...they really cultivate the young alien mind.

OpenStudy (mr.math):

Sometimes the thought that he has got supernatural powers occurs to me :-P

OpenStudy (asnaseer):

:D lol! Zarkon

OpenStudy (anonymous):

Myin, more than they could possibly spend, (they spend most of the time doing maths)

OpenStudy (anonymous):

one time someone asked me why they call it "telescope" which was the first time i realized that it was not obvious. i guess modern telescopes don't actually "telescope"

myininaya (myininaya):

Zarkon can you use that money to see me to plant doom university.

OpenStudy (mr.math):

Me too please!

myininaya (myininaya):

send* lol

OpenStudy (anonymous):

@send?

OpenStudy (zarkon):

only if you lived there...the out of planet tuition is outragious.

OpenStudy (asnaseer):

I'd be happy to stow away in the luggage compartment!

myininaya (myininaya):

do you fold well?

OpenStudy (mr.math):

haha

OpenStudy (asnaseer):

@satellite - we still have telescopic aerials, poles, etc - you could use those as examples? :) @myininaya - yes - I was trained especially for that by being fed lots of soft bananas :D

OpenStudy (mr.math):

Looks like asnaseer has been to other planets too, haven't you?

OpenStudy (asnaseer):

Shhhh.... Mr.Math - you are ruining my plan to rule the Universe.... ;-)

myininaya (myininaya):

Well anyways great question and great answer to all of the aliens involved.

OpenStudy (anonymous):

Myin: Take a look at this M.SE thread ( http://math.stackexchange.com/questions/18576/)

OpenStudy (mr.math):

Whops! sorry.

myininaya (myininaya):

so this is where you got the problem ffm? from this universe?

OpenStudy (anonymous):

lol no, I just got it from MR.M's previous question.